# Understanding and utilizing the inner bonds of process tensors

### v1

@@ II.1 Notation
$o_{(\nu,\mu),(\xi,\eta)}=$ $\big{(}\langle\mu|\otimes\langle\xi|\big{)}\hat{O}\big{(}|\nu\rangle\otimes|\eta\rangle\big{)},$ (8)
@@ III.2 MPO compression
where $U$ and $V$ are matrices with orthogonal columns, while $\Sigma$ is a diagonal matrix with the non-negative real singular values $\sigma_{j}$ on the diagonal, which are assumed to be sorted in descending order, i.e. $\sigma_{j}\geq\sigma_{k}$ for $j\leq k$. Here, we use truncated SVDs by keeping only the $s$ most significant singular values $\sigma_{j}\geq\epsilon\sigma_{0}$, where $\epsilon$ is a given truncation threshold relative to the largest singular value $\sigma_{0}$, and also restricting the matrices $U$ and $V$ to their first $s$ columns. For our applications, $s$ will be the inner bond dimension after PT-MPO compression. In fact, the Eckart-Young-Mirsky theorem [51] states that SVDs provide the best low-rank approximation to a general matrix for given rank $s$. To simplify the notation, we henceforth imply truncation whenever we refer to SVDs like in Eq. (25).
@@ IV.2 Mixed system-environment observables: Currents
To infer physical insights from measurement of currents through zero-dimensional quantum structures like single molecules or quantum dots, a comparison with theoretical predictions is highly desirable [53]. A typical scenario for charge transport involves a single quantum site, which is described by a two-level system (the site being either occupied or not), and which is coupled to two metallic leads at different chemical potentials $\mu_{1}>\mu_{2}$. The leads are modelled as fermionic baths. The total Hamiltonian is $H=H_{S}+H_{E_{1}}+H_{E_{2}}$ with
$H_{E_{i}}=$ $\sum_{k}\hbar\omega_{i,k}c^{\dagger}_{i,k}c_{i,k}+\sum_{k}\hbar g_{i,k}\big{(}c^{\dagger}_{i,k}\sigma^{-}+c_{i,k}\sigma^{+}\big{)},$ (46)
where $c^{(\dagger)}_{i,k}$ denote annihilation (creation) operators for an excitation in the $k$-th mode of the $i$-th environment, $\hbar\omega_{i,k}$ and $g_{i,k}$ are the corresponding energies and coupling constants, and $\sigma^{+}$ and $\sigma^{-}$ describe the raising and lowering of population on the central site. We set $H_{S}=0$, and assume that the two-level system is initially unoccupied, the first lead is completely filled and the second lead is completely empty (chemical potentials $\mu_{1/2)}=\pm\infty$). The energies and couplings for both baths are obtained by uniformly discretizing the spectral densities $J_{i}(\omega)=\sum_{k}g_{i,k}^{2}\delta(\omega-\omega_{i,k})$ of the shape of a bump function
@@
$=$ $2\sum_{k}g_{i,k}\textrm{Im}\big{\{}\langle c^{\dagger}_{i,k}\sigma^{-}\rangle\big{\}}.$ (48)
Hence, the current $I_{i}$ can be extracted from the inner bonds of the PT-MPO describing the influence of the bath $E_{i}$ by tracking how the mixed system-environment operator $\sum_{k}g_{i,k}c^{\dagger}_{i,k}\sigma^{-}$ transforms under MPO compression.
@@
Moreover, two PT-MPOs describing the two leads can be calculated independently of each other and combined when propagating the two-level system dynamics (see Refs. [17, 37] for applications). The fact that this treatment remains numerically exact can be seen straightforwardly in the notation developed here: Using transformation matrices $\mathcal{T}^{[i]}$ and PT-MPO matrices $\mathcal{Q}^{[i]}=\mathcal{T}^{[i]}e^{\mathcal{L}_{E,i}}\big{(}\mathcal{T}^{[i]}\big{)}^{-1}$ for bath $i$, we find that
$\mathcal{Q}^{[1]}\mathcal{Q}^{[2]}=$ $\mathcal{T}^{[1]}\mathcal{T}^{[2]}e^{\mathcal{L}_{E,1}\Delta t}e^{\mathcal{L}_{E,2}\Delta t}\big{(}\mathcal{T}^{[2]}\big{)}^{-1}\big{(}\mathcal{T}^{[1]}\big{)}^{-1}$
$=$ $\mathcal{T}e^{(\mathcal{L}_{E,1}+\mathcal{L}_{E,2})\Delta t}\mathcal{T}^{-1}+\mathcal{O}(\Delta t^{2})$ (49)
@@
Figure 3: Occupations (a) and currents (b) in charge transport between a single two-level system and one or two metallic leads at chemical potentials $\mu_{1/2}=\pm\infty$, as depicted in the inset of (b). The simulation for open leads includes Lindblad terms for states in the leads that restore their initial occupation with a rate equal to $\kappa$ (see main text). Dashed black lines depict results for the single lead case in the Markov limit.
@@
In Fig. 3(a) and (b) we show the two-level system occupations and the particle current from the system to the first lead, respectively, for simulations involving either only the first lead or both leads. The PT-MPOs are calculated using parameters $\omega_{BW}=64\kappa$, $N=128$, $\Delta t=0.05/\kappa$, and $\epsilon=10^{-7}$.
@@
In the case of a single lead with large chemical potential, carriers flow from lead to the two-level system (negative $I_{1}$) until the latter is fully populated. The dynamics is largely in agreement with the analytical behavior $n_{S}(t)=1-e^{-\kappa t}$ in the Markov limit, which is expected to hold as the bath correlation time $\sim 1/\omega_{BW}$ is short compared to the relaxation time $1/\kappa$. Due to the conservation of particle number and the coupling to only a single lead, the current can be obtained from the system evolution via $I_{1}(t)=-\frac{\partial}{\partial}n_{S}(t)$, which is $I_{1}(t)=-\kappa e^{-\kappa t}$ in the Markov limit.
For a site coupled to two leads, changes of site occupations are due to currents to either lead, which obfuscates the relation between currents through one of the leads and system observables. This is where extracting the current via the mixed system-environment observable defined in Eq. (48) is useful. In this two-lead situation, we find that the site occupation is nearly constant after a time $\sim 1/\kappa$. The current remains non-zero, which indicates non-vanishing steady-state current from the lead with high chemical potential to the lead with low chemical potential. However, the current does not become constant within the total propagation time in Fig. 3. This is explained by the fact that the PT-MPOs describe the coupling between site and leads to all orders, which includes phase space filling effects: Due to the charge transport from the initially fully occupied lead to the initially empty lead, the former becomes depleted while the latter becomes filled. This eventually reduces the magnitude of the current.
This phenomenon indicates that the model used so far lacks physical effects like thermalization within the leads as well as the power source connected to them, which keep the chemical potential at a fixed value. This can be modelled within an open quantum systems approach, where the environment Liouvillians additionally contain Lindblad terms that eventually restore the initial distributions of lead state occupations
$\mathcal{L}_{E,1}=$ $-\frac{i}{\hbar}[H_{E_{1}},\,.\,]+\gamma_{1}\sum_{k}\bigg{[}c^{\dagger}_{1,k}\,.\,c_{1,k}-\frac{1}{2}\big{\{}c_{1,k}c^{\dagger}_{1,k},\,.\,\big{\}}\bigg{]},$ (50a)
$\mathcal{L}_{E,2}=$ $-\frac{i}{\hbar}[H_{E_{2}},\,.\,]+\gamma_{2}\sum_{k}\bigg{[}c_{2,k}\,.\,c^{\dagger}_{2,k}-\frac{1}{2}\big{\{}c^{\dagger}_{2,k}c_{2,k},\,.\,\big{\}}\bigg{]}.$ (50b)
The corresponding results for a site coupled to two open leads is also shown in Fig. 3, where we set $\gamma_{1}=\gamma_{2}=\kappa$. There, the current indeed becomes essentially stationary. Note, however, that the exact value depends on the rates $\gamma_{i}$, so these should be considered true physical parameters. Finally, the small kinks in the currents towards the end of the propagation time are related to convergence as discussed in the next example.
@@ IV.4 Inner bonds of PT-MPOs from different algorithms: Example in quantum thermodynamics
We apply this approach to a test case in quantum thermodynamics [54, 55]. In this research area, one investigates how concepts of macroscopic thermodynamics can be generalized to quantum mechanical systems. While work and heat play a central role in classical thermodynamics, their definition is more complex in quantum settings [54]. Irrespective of such subtleties we are here interested in calculating the total energy absorbed by a quantum system subject to external driving. If this system is a non-Markovian open quantum system, a further challenge is to resolve how the total absorbed energy is distributed over different terms in the Hamiltonian, namely the mean system energy, the mean energy absorbed into purely environmental degrees of freedom, and the mean system-environment interaction energy, which is a consequence of non-vanishing system-environment correlations. This information can be readily obtained from PT-MPOs via their inner bonds.
@@
We are interested in the energetics, i.e., how the energy is distributed over time between the different terms in the total Hamiltonian. $\langle H_{S}(t)\rangle$ as well as $\langle H_{PS}(t)\rangle$ can be directly obtained from the reduced system density matrix. Because system observables are directly driven by the interaction term $H_{I}$, the observable closure for the environment observable $\hat{O}_{1}=\sum_{k}\hbar g_{k}b^{\dagger}_{k}$ converges quickly ($\sum_{k}\hbar g_{k}b_{k}=\hat{O}_{1}^{\dagger}$ does not have to be calculated separately). Thus, we extract the mean interaction energy $\langle H_{I}(t)\rangle$ via the correspondig closure. Finally, the free phonon energy enters the equation of motion for the system observables only indirectly by affecting system-bath correlations. As discussed in the previous example, this implies that faster convergence is expected when the increase of the free phonon energy with respect to its initial value $\langle\Delta H_{E}^{0}(t)\rangle=\langle H_{E}^{0}(t)\rangle-\langle H_{E}^{0}(0)\rangle$ is not extracted directly but instead by integrating the equation of motion

### v2

@@ II.1 Notation
$o_{(\nu,\mu),(\xi,\eta)}=$ $\big{(}\langle\mu|\otimes\langle\eta|\big{)}\hat{O}\big{(}|\nu\rangle\otimes|\xi\rangle\big{)},$ (8)
@@ III.2 MPO compression
where $U$ and $V$ are matrices with orthogonal columns, while $\Sigma$ is a diagonal matrix with the non-negative real singular values $\sigma_{j}$ on the diagonal, which are assumed to be sorted in descending order, i.e. $\sigma_{j}\geq\sigma_{k}$ for $j\leq k$. Here, we use truncated SVDs by keeping only the $s$ most significant singular values $\sigma_{j}\geq\epsilon\sigma_{0}$, where $\epsilon$ is a given truncation threshold relative to the largest singular value $\sigma_{0}$, and also restricting the matrices $U$ and $V$ to their first $s$ columns. For our applications, $s$ will be the inner bond dimension after PT-MPO compression.
In fact, the Eckart-Young-Mirsky theorem [51] states that SVDs provide the best low-rank approximation to a general matrix for given rank $s$. However, MPO compression at a given site using truncated SVD is only locally optimal if the MPO is in mixed canonical form [21]. One way to achieve this is to first perform a forward sweep with non-truncating SVDs and only truncating during the backward sweep. Yet local optimal compression is not necessarily needed for the most efficient algorithm. Here, we use a strategy in the spirit of the zip-up algorithm of Ref. [52], where truncation is performed for both, forward and backward sweeps. The sizable speed-up achieved by operating on matrices with smaller dimensions is typically well worth the slight increase in numerical error, especially because the latter can be mitigated by using a smaller compression threshold. As discussed in Appendix A of Ref. [29] as well as in Ref. [53], choosing different compression parameters for different sweeps can be used to fine-tune PT-MPO algorithms. For example, using a different threshold $\epsilon_{fw}$ for forward sweeps compared to backward sweeps $\epsilon_{bw}$, one can interpolate between the locally optimal compression ($\epsilon_{fw}=0$) and the zip-up-like approach ($\epsilon_{fw}=\epsilon_{bw}$). Here, we use $\epsilon_{fw}=\epsilon_{bw}=\epsilon$ throughout this article. Furthermore, note that MPO compression by SVDs in general is aimed at minimizing the norm distance between MPOs before and after compression. To our knowledge, no formal theory linking this norm distance to a precise error bound for physical observables is available so far. However, numerical simulations in Ref. [53] indicate that the PT-MPO norm distance and system observables show similar convergence behavior. This provides empirical justification for PT-MPO compression using sweeps with SVDs as described above.
Next, we summarize how MPO matrices and in particular their inner bonds are affected by forward and backward sweeps. To simplify the notation, we henceforth imply truncation whenever we refer to SVDs like in Eq. (25).
@@ IV.2 Mixed system-environment observables: Currents
To infer physical insights from measurement of currents through zero-dimensional quantum structures like single molecules or quantum dots, a comparison with theoretical predictions is highly desirable [55]. A typical scenario for charge transport involves a single quantum site, which is described by a two-level system (the site being either occupied or not), and which is coupled to two metallic leads at different chemical potentials $\mu_{1}>\mu_{2}$. The leads are modelled as fermionic baths. The total Hamiltonian is $H=H_{S}+H_{E_{1}}+H_{E_{2}}$ with $H_{E_{i}}=\sum_{k}H_{E_{i}}^{(k)}$ and
$H_{E_{i}}^{(k)}=$ $\hbar\omega_{i,k}c^{\dagger}_{i,k}c_{i,k}+\hbar g_{i,k}\big{(}c^{\dagger}_{i,k}c_{S}+c^{\dagger}_{S}c_{i,k}\big{)},$ (46)
where $c^{(\dagger)}_{i,k}$ denote fermionic annihilation (creation) operators for the $k$-th mode of the $i$-th environment, $\hbar\omega_{i,k}$ and $g_{i,k}$ are the corresponding energies and coupling constants, and $c^{\dagger}_{S}$ and $c_{S}$ describe fermionic operators of the system mode. We set $H_{S}=0$, and assume that the central site is initially unoccupied, the first lead is completely filled and the second lead is completely empty (chemical potentials $\mu_{1/2}=\pm\infty$). The energies and couplings for both baths are obtained by uniformly discretizing the spectral densities $J_{i}(\omega)=\sum_{k}g_{i,k}^{2}\delta(\omega-\omega_{i,k})$ of the shape of a bump function
@@
$=$ $2\sum_{k}g_{i,k}\textrm{Im}\big{\{}\langle c^{\dagger}_{i,k}c_{S}\rangle\big{\}}.$ (48)
This is a mixed system-environment observable, which depends on correlations between the system and the environment. At first glance, this observable as well as the environment Hamiltonian in Eq. (46) seem decomposable into sums of terms each involving only creation and annihilation operators of a single environment mode, as we require for the environment observable extraction described in this article. Note, however, that fermionic operators obey the canonical anticommutator relations, whereas the decomposition into independent modes so far implicitly assumed that operators for different modes commute. As described in more detail in Ref. [53], PT-MPOs obeying the proper fermionic anticommutator relations can be obtained by a slight modification of ACE based on a Jordan-Wigner transformation. Here, we briefly summarize the main aspects while using a formulation that facilitates the extraction of the current $I_{1}(t)$ in a setup with up to two environments.
First, the Jordan-Wigner transformations requires an ordering of fermionic modes. Here, we assume the order
$(1,1),(1,2),\dots,(1,N_{E_{1}}),(S),(2,N_{E_{2}}),\dots,(2,2),(2,1),$ (49)
where (S) denotes the system mode and $(i,k)$ denotes the $k$-th mode of the $i$-th environment. The anticommutation relations remain fulfilled when the fermionic operators are replace by spin-1/2 climbing operators
$c_{i,k}=$ $P^{[(1,1),(i,k-1)]}\sigma^{-}_{i,k},$ (50a)
$c^{\dagger}_{i,k}=$ $P^{[(1,1),(i,k-1)]}\sigma^{+}_{i,k},$ (50b)
where
$P^{[(i_{1},k_{1}),(i_{2},k_{2})]}=$ $\prod_{(i,k)=(i_{1},k_{1})}^{(i_{2},k_{2})}(-\sigma^{z}_{i,k})$ (51)
is the product of parity operators from mode $(i_{1},k_{1})$ to $(i_{2},k_{2})$ in the order given by Eq. (49), which takes the value +1 (-1) if the modes in the range have an even (odd) number of occupations. Applying this transformation to the Hamiltonian in Eq. (46), one can show that
$H_{E_{i}}^{(k)}$ $=\big{(}P^{[(1,1),(i,k-1)]}_{1}\big{)}^{\sigma^{+}_{S}\sigma^{-}_{S}}\tilde{H}_{E_{i}}^{(k)}\big{(}P^{[(1,1),(i,k-1)]}_{1}\big{)}^{\sigma^{+}_{S}\sigma^{-}_{S}},$ (52)
where
$\tilde{H}_{E_{i}}^{(k)}=$ $\hbar\omega_{i,k}\sigma^{+}_{i,k}\sigma^{-}_{i,k}+\hbar g_{i,k}\big{(}\sigma^{+}_{i,k}\sigma^{-}_{S}+\sigma^{+}_{S}\sigma^{-}_{i,k}\big{)}$ (53)
is the spin analogue of the fermionic Hamiltonian $H_{E_{i}}^{(k)}$.
Now, while the fermionic mode Hamiltonians individually contain non-local terms via the parity operators, it can be shown that these non-local terms cancel when the propagators of all modes are combined, leaving only local parity terms [53]. Specifically, representing the PT-MPO matrices for environment $i$ using lossy compression matrices as well as the symmetric Trotter decomposition of the full environment propagator in Eq. (19), one finds
$\mathcal{Q}^{[i]}$
$=\mathcal{T}e^{\mathcal{L}_{E_{i}}^{(N_{E_{i}})}\frac{\Delta t}{2}}\dots e^{\mathcal{L}_{E_{i}}^{(2)}\frac{\Delta t}{2}}e^{\mathcal{L}_{E_{i}}^{(1)}\Delta t}e^{\mathcal{L}_{E_{i}}^{(2)}\frac{\Delta t}{2}}\dots e^{\mathcal{L}_{E_{i}}^{(N_{E_{i}})}\frac{\Delta t}{2}}\mathcal{T}^{-1}$
$=\mathcal{T}\mathcal{B}^{[i,N_{E_{i}}]}\dots\mathcal{B}^{[i,2]}\mathcal{B}^{[i,1]}\mathcal{B}^{[i,1]}\mathcal{B}^{[i,2]}\dots\mathcal{B}^{[i,N_{E_{i}}]}\mathcal{T}^{-1},$ (54)
where
$\mathcal{B}^{[i,k]}:=$ $e^{\tilde{\mathcal{L}}_{E_{i}}^{(k)}\frac{\Delta t}{2}}\Big{[}\big{(}-\sigma^{z}_{i,k}\big{)}^{\sigma^{+}_{S}\sigma^{-}_{S}}\otimes\big{(}-\sigma^{z}_{i,k}\big{)}^{\sigma^{+}_{S}\sigma^{-}_{S}}\Big{]}$ (55)
are the local environment mode propagators corresponding to the spin Hamiltonians $\tilde{H}_{E_{i}}^{(k)}$ in Eq. (53) modified by local parity operators $\big{(}-\sigma^{z}_{i,k}\big{)}^{\sigma^{+}_{S}\sigma^{-}_{S}}$. To summarize, PT-MPOs correctly accounting for fermionic anticommutation can be obtained simply by replacing the individual mode propagators $\mathcal{B}$ in Eq. (22) by corresponding $\mathcal{B}^{[i,k]}$ in Eq. (55).
To extract the particle current via the inner bonds of the fermionic PT-MPO, we apply the Jordan-Wigner transform to Eq. (48), which yields
$I_{i}:=2\sum_{k}g_{1,k}\textrm{Im}\big{\{}\langle\sigma^{+}_{i,k}P^{[k+1,N_{E_{i}}]}_{1}\sigma^{-}_{S}\rangle\big{\}}.$ (56)
Again the non-local parity terms make adaptations necessary. This is achieved by replacing the iteration in Eq. (43) by
$\mathfrak{o}_{d_{l}^{(k)}}=$ $\sum_{d_{l}^{(k-1)}}\sum_{\beta_{l}^{(k)},\xi,\eta}\delta_{\beta_{l}^{(k)},(\xi,\eta)}\Big{(}\mathfrak{o}_{d_{l}^{(k-1)}}\langle\eta|-\sigma^{z}|\xi\rangle$
$+q_{d_{l}^{(k-1)}}\langle\eta|\sigma^{+}|\xi\rangle\Big{)}\overleftrightarrow{T}^{-1}_{(d_{l}^{(k-1)},\beta_{l}^{(k)}),d_{l}^{(k)}}.$ (57)
With this observable closure, we are now in the position to extract the fermionic particle current $I_{i}(t)$ as a mixed system-environment observable from the inner bonds of the fermionic PT-MPO.
@@
Finally, we point out that the numerically exact simulation in the presence of two generally non-Markovian environments is possible by propagating the time evolution with two PT-MPOs, which are calculated independently of each other. This was utilized already for bosonic environments, e.g., in Refs. [17, 37]. Here, the ordering used for the Jordan-Wigner transformation in Eq. (49) ensures that no additional non-local parity operators coupling the two environments emerge. The fact that this treatment remains numerically exact can be seen straightforwardly in the notation developed here: Using transformation matrices $\mathcal{T}^{[i]}$ and PT-MPO matrices $\mathcal{Q}^{[i]}=\mathcal{T}^{[i]}e^{\mathcal{L}_{E_{i}}\Delta t}\big{(}\mathcal{T}^{[i]}\big{)}^{-1}$ for bath $i$, we find that
$\mathcal{Q}^{[1]}\mathcal{Q}^{[2]}=$ $\mathcal{T}^{[1]}\mathcal{T}^{[2]}e^{\mathcal{L}_{E_{1}}\Delta t}e^{\mathcal{L}_{E_{2}}\Delta t}\big{(}\mathcal{T}^{[2]}\big{)}^{-1}\big{(}\mathcal{T}^{[1]}\big{)}^{-1}$
$=$ $\mathcal{T}e^{(\mathcal{L}_{E_{1}}+\mathcal{L}_{E_{2}})\Delta t}\mathcal{T}^{-1}+\mathcal{O}(\Delta t^{2})$ (58)
@@
Figure 3: Occupations (a) and currents (b) in charge transport between a single site and one or two metallic leads at chemical potentials $\mu_{1/2}=\pm\infty$, as depicted in the inset of (b). Dashed and dotted black lines depict results in the Markov limit for the single lead case and the two lead case, respectively.
@@
In Fig. 3(a) and (b) we show the system site occupations and the particle current from the system to the first lead, respectively, for simulations involving either only the first lead or both leads. The PT-MPOs are calculated using parameters $\omega_{BW}=64\kappa$, $N=128$, $\Delta t=0.025/\kappa$, and $\epsilon=10^{-7}$.
@@
In the case of a single lead with large chemical potential, carriers flow from lead to the system site (negative $I_{1}$) until the latter is fully populated. The dynamics is largely in agreement with the analytical behavior $n_{S}(t)=1-e^{-\kappa t}$ in the Markov limit, which is expected to hold as the bath correlation time $\sim 1/\omega_{BW}$ is short compared to the relaxation time $1/\kappa$. Due to the conservation of particle number and the coupling to only a single lead, the current can be obtained from the system evolution via $I_{1}(t)=-\frac{\partial}{\partial}n_{S}(t)$, which is $I_{1}(t)=-\kappa e^{-\kappa t}$ in the Markov limit.
For a site coupled to two leads, changes of site occupations are due to currents to either lead, which generally obfuscates the relation between currents through one of the leads and system observables. This is where extracting the current via the mixed system-environment observable defined in Eq. (48) is useful. In the present scenario where initially one lead is fully occupied and one lead is completely empty, the dynamics can be compared to its Markovian limit, where it is governed by the rate equation $\frac{\partial}{\partial t}n_{S}=-I_{1}(t)-I_{2}(t)$, where $I_{1}(t)=\kappa(n_{S}(t)-1)$ and $I_{2}(t)=\kappa n_{S}(t)$ are the Markovian currents to environment 1 and 2, respectively. This equation of motion is solved by system occupations $n_{S}(t)=\frac{1}{2}(1-e^{-2\kappa t})$, from which we get the current $I_{1}(t)=-\frac{\kappa}{2}(1+e^{-2\kappa t})$. As can be seen in Fig. 3, for the case of two leads, simulated site occupations as well as currents also match the Markovian prediction well after a short initial phase on the timescale $\sim 1/\omega_{BW}$.
@@ IV.4 Inner bonds of PT-MPOs from different algorithms: Example in quantum thermodynamics
We apply this approach to a test case in quantum thermodynamics [56, 57, 58, 59, 60]. In this research area, one investigates how concepts of macroscopic thermodynamics can be generalized to quantum mechanical systems. While work and heat play a central role in classical thermodynamics, their definition is more complex in quantum settings [61, 62]. Irrespective of such subtleties we are here interested in calculating the total energy absorbed by a quantum system subject to external driving. If this system is a non-Markovian open quantum system, a further challenge is to resolve how the total absorbed energy is distributed over different terms in the Hamiltonian, namely the mean system energy, the mean energy absorbed into purely environmental degrees of freedom, and the mean system-environment interaction energy. The latter is a consequence of non-vanishing system-environment correlations and highlights the need for methods in quantum thermodynamics that can account for strong system-environment coupling [56, 57, 58, 59, 60]. Here, the energy distribution over the different terms can be readily obtained from PT-MPOs via their inner bonds.
@@
We are interested in the energetics, i.e., how the energy is distributed over time between the different terms in the total Hamiltonian. $\langle H_{S}(t)\rangle$ as well as $\langle H_{PS}(t)\rangle$ can be directly obtained from the reduced system density matrix. Because system observables are directly driven by the interaction term $H_{I}$, the observable closure for the environment observable $\hat{O}_{1}=\sum_{k}\hbar g_{k}b^{\dagger}_{k}$ converges quickly ($\sum_{k}\hbar g_{k}b_{k}=\hat{O}_{1}^{\dagger}$ does not have to be calculated separately). Thus, we extract the mean interaction energy $\langle H_{I}(t)\rangle$ via the corresponding closure. Finally, the free phonon energy enters the equation of motion for the system observables only indirectly by affecting system-bath correlations. As discussed in the previous example, this implies that faster convergence is expected when the increase of the free phonon energy with respect to its initial value $\langle\Delta H_{E}^{0}(t)\rangle=\langle H_{E}^{0}(t)\rangle-\langle H_{E}^{0}(0)\rangle$ is not extracted directly but instead by integrating the equation of motion